OpenStudy (anonymous):

help>???? 55x + 35x = 2

OpenStudy (anonymous):

your maths is very poor

OpenStudy (uri):

90x=2 x=90/2 Divide and ans.

OpenStudy (phi):

if you have 55 x's and add 35 more x's, how many x's do you have ?

OpenStudy (uri):

@bhusal This site is to help.

OpenStudy (anonymous):

i think its 45

OpenStudy (anonymous):

x=2/90=1/45

OpenStudy (uri):

@katt97 Stop giving direct answers.

OpenStudy (anonymous):

ya this site is to help but should self try and then and only ask

OpenStudy (phi):

you should get 90 x= 2 now divide both sides by 90, like this x * 90/90 = 2/90 that simplifies to x = 1/45

OpenStudy (anonymous):

OKAY, so 55 + 35 because its x then divide by 2? give 45.

OpenStudy (anonymous):

@bhusal um, excuse me I am from spain and it is really hard coming to america without math skills and please don't make fun of me I do try. :/

OpenStudy (phi):

55x + 35x means you have 90 x's, written 90x so 55x + 35x = 2 is the same as 90 x = 2 to find x, divide both sides by 90

OpenStudy (anonymous):

OKAY, thank y ou so much!!!!! :)

OpenStudy (anonymous):

I think i am getting better and better. :) ty...!!!

OpenStudy (phi):

the answer is not 45. it is 2/90 which can be simplified to 1/45

OpenStudy (anonymous):

oh i am very sorry but u need to work hard and i am not trying to make fun of you please dont mind think positive and work had. have ego

OpenStudy (anonymous):

i am working hard ??? :) ty.